a determinant is chosen at random from the set of all determinant of order 2 with element elements 0 or 1
only.Find the probabilty that the determinant chosen is non zero
 

Grade:11

1 Answers

ande yashwanth
7 years ago
There are four elements in the determinant​.Each element can be taken in two ways (0 or 1). So total possible determinants are 2^4.In this if the principal diagonal elements are 1 and a12,a21 elements are 1 then the determinant is non zero.(only two cases)Probability=2/2^4=1/8